When pond water is examined under the microscope there are nasty bugs to be seen. The average concentration of these bugs is 3 per milliliter. Determine the probability that a random sample of 5 milliliters of pond water contains exactly 14 bugs:
a) Using the Poisson distribution.
b) Using the Normal approximate to the Poisson distribution.
1
Expert's answer
2020-11-26T17:51:13-0500

The provided mean is 

"\lambda= 3"


a)We need to compute Pr(X=5). Therefore, the following is obtained using ti84,Pr(X=5)=0.10082


b)Using the Normal approximate to the Poisson distribution.

mean = "n*\lambda" =5*3=15

sd= "\sqrt{(n*\lambda)}" ="\sqrt{15}=" 3.873


Hence P(x=14)= =NORMSDIST((14.5-15)/SQRT(15))-NORMSDIST((13.5-15)/SQRT(15))=0.099372